# Does hot QCD have a conformal manifold in the chiral limit?

### 10 Mar 2026 v1

@@ Abstract
The PCB resummation then constructs Padé approximants to $\overline{\mathcal{F}}(z)$, instead of to $\mathcal{F}(\delta)$. An $[m,n]$ Padé approximant to $\overline{\mathcal{F}}(z)$ is a rational function of the form $P_{m}(z)/Q_{n}(z)$ that matches the series expansion of $\overline{\mathcal{F}}(z)$ to order $z^{m+n}$. The $[1,1]$, $[2,1]$ and $[1,2]$ Padé approximants are then
$\overline{\mathcal{F}}(z)\big|_{[1,1]}$ $=\frac{1-2z}{2(p-1)z+1}\,$, (S4.10)
$\overline{\mathcal{F}}(z)\big|_{[2,1]}$ $=\frac{-2pz\left(4p^{2}z-2p+z-2\right)-2p-3z+2}{p(8z-2)-3z+2}\,$, (S4.11)
$\overline{\mathcal{F}}(z)\big|_{[1,2]}$ $=\frac{\left(4p^{2}-3\right)z+2}{z\left(2p\left(4p^{2}z+2p+z+2\right)-3\right)+2}\,.$ (S4.12)
When $p$ takes values in an appropriate range, these Padé approximants do not put spurious poles on the interval $z\in[0,1]$, and are then reasonable to use in our resummation. This range of values is $p\in(1/2,\infty)$ for the $[1,1]$ approximant, $p\in(1/6,1]$ for the $[2,1]$ approximant, and $p\in(p_{*},\infty)$ where $p_{*}\approx 0.1478$ is the real root of $8p^{3}+4p^{2}+6p-1$ for the $[1,2]$ approximant.
The final step is to perform the Borel integral to $\mathcal{F}(\delta)|_{[m,n]}=\overline{\mathcal{F}}(z(\delta))|_{[m,n]}$. We thus obtain PCB-resummed approximants of $F(\epsilon)$, controlled by parameter $p$:
@@
$F(\epsilon)\big|_{[m,n]}=\frac{1}{\epsilon}\int_{0}^{\infty}d\delta\ e^{-\delta/\epsilon}\mathcal{F}(\delta)\big|_{[m,n]}\,.$ (S4.13)
We summarize our results at $\epsilon=1$ in Table 3, where the error bars correspond to the envelope of $F(1)$ as $p$ varies over its allowed range.
$[m,n]$ $[1,1]$ $[2,1]$ $[1,2]$
$F(1)|_{[m,n]}$ $0.70\pm 0.03$ $0.70\pm 0.02$ $0.71\pm 0.03$
Table 3: Estimates for $F(\epsilon)$ at $\epsilon=1$ obtained via Padé-Conformal-Borel (PCB) resummation of the three-term truncated small-$\epsilon$ asymptotic series using $[1,1]$, $[2,1]$, and $[1,2]$ Padé approximants. For each approximant, the quoted uncertainty is the envelope of results under variation of the assumed location p-p of the leading Borel-plane singularity, over the range for which the Borel-Laplace integral is well-defined (no poles on the integration contour).
As a conservative overall estimate for $F(1)$, we take the union of the three envelopes, yielding $F(1)=0.71\pm 0.04$. We show the resulting estimates for $\Delta_{\phi}(1)$ for several values of $N_{f}$ in Table 4. We note that the results do satisfy the unitarity bound $\Delta_{\phi}\geq\frac{1}{2}$ in 3D.
$N_{f}$ 22 33 44 55 66 77 88
$\Delta_{\phi}$ $0.53\pm 0.03$ $0.63\pm 0.03$ $0.67\pm 0.04$ $0.68\pm 0.04$ $0.69\pm 0.04$ $0.70\pm 0.04$ $0.70\pm 0.04$
@@
Table 4: Estimates for $\Delta_{\phi}$ in 33D obtained from the Padé-Conformal-Borel (PCB) resummation of the first three terms in a $(2+\epsilon)$D expansion as a function of $N_{f}$.

### 14 Apr 2026 v2

@@ Abstract
An $[m/n]$ Padé approximant to $\mathcal{B}[\Delta_{E}](\delta)$ is a degree-$(m,n)$ rational function that matches the series expansion of $\mathcal{B}[\Delta_{E}](\delta)$ to order $\delta^{m+n}$. The near-diagonal Padé approximants at order $\delta^{4}$ are
$\mathcal{B}[\Delta_{E}](\delta)\big|_{[3/1]}$ $=\frac{2-\Bigl(\frac{5}{8}\!+\!\frac{9\zeta_{3}}{2N_{f}^{2}}\Bigr)\delta-\frac{1}{4}\delta^{2}+\Bigl(\frac{7}{192}\!+\!\frac{9\zeta_{3}}{16N_{f}^{2}}\Bigr)\delta^{3}}{1-\Bigl(\frac{5}{16}\!+\!\frac{9\zeta_{3}}{4N_{f}^{2}}\Bigr)\delta}\,$, (S4.7a)
$\mathcal{B}[\Delta_{E}](\delta)\big|_{[2/2]}$ $=\frac{2-\frac{1}{3}\delta-\Bigl(\frac{43}{144}\!+\!\frac{3\zeta_{3}}{4N_{f}^{2}}\Bigr)\delta^{2}}{1-\frac{1}{6}\delta-\Bigl(\frac{7}{288}\!+\!\frac{3\zeta_{3}}{8N_{f}^{2}}\Bigr)\delta^{2}}\,$, (S4.7b)
$\mathcal{B}[\Delta_{E}](\delta)\big|_{[1/3]}$ $=\frac{2-\Bigl(\frac{17}{8}\!+\!\frac{9\zeta_{3}}{2N_{f}^{2}}\Bigr)\delta}{1-\Bigl(\frac{17}{16}\!+\!\frac{9\zeta_{3}}{4N_{f}^{2}}\Bigr)\delta+\frac{1}{8}\delta^{2}-\Bigl(\frac{43}{384}\!+\!\frac{9\zeta_{3}}{32N_{f}^{2}}\Bigr)\delta^{3}}\,.$ (S4.7c)
All of them have poles on the positive real axis. We then obtain the $[m/n]$ PB-resummed function $\Delta_{E}(\epsilon)$ by performing the Borel integral,
@@
$\Delta_{E}(\epsilon)\big|_{[m/n]}=\ \lim_{\theta\to 0}\ \frac{1}{\epsilon}\int_{0}^{\infty}d\delta\,\exp\left(-\delta e^{i\theta}/\epsilon\right)\,\mathcal{B}[\Delta_{E}](\delta e^{i\theta})\big|_{[m/n]}\,.$ (S4.8)
In a resurgent transseries representation of $\Delta_{E}(\epsilon)$, $\mathop{\mathrm{im}}\Delta_{E}(\epsilon)\big|_{[m/n]}$ would be canceled by a corresponding imaginary part of a non-perturbative term in the transseries, which would also give a real contribution of approximately the same magnitude. We thus use the magnitude of the imaginary ambiguities in the Padé-Borel resummations as an estimate of the error on our extrapolations.
We show the functions $\Delta_{E}(\epsilon)|_{[m/n]}$ with $N_{f}=2\sim 7$ in Fig. 2. We summarize our $\epsilon=1$ results as a function of $[m/n]$ and $N_{f}$ in Table 3. We use the union of the envelopes of $\Delta_{E}(1)|_{[3/1]}$, $\Delta_{E}(1)|_{[2/2]}$, and $\Delta_{E}(1)|_{[1/3]}$ to determine our estimation of 3D $\Delta_{E}$ and the extrapolation error, as we summarize in Table 4. The extrapolation errors are large with the currently available four-loop data, and thus higher-loop data would be very valuable.
Figure 2: Evaluations of the PB-resummed function $\Delta_{E}(\epsilon)$ with $N_{f}=2\sim 7$. The solid lines show $\mathrm{Re}\,\Delta_{E}(\epsilon)$ while the shadow bands indicate error bands determined from the imaginary ambiguities in PB resummation.
$[m/n]$ $[3/1]$ $[2/2]$ $[1/3]$
$\Delta_{E}(1),\ {N_{f}=2}$ $1.64\pm\,0.05$ $1.94\pm\,0.56$ $1.76\pm\,0.01$
$\Delta_{E}(1),\ {N_{f}=3}$ $1.71\pm\,0.18$ $1.76\pm\,0.67$ $1.77\pm\,0.01$
$\Delta_{E}(1),\ {N_{f}=4}$ $1.72\pm\,0.31$ $1.64\pm\,0.71$ $1.77\pm\,0.02$
$\Delta_{E}(1),\ {N_{f}=5}$ $1.69\pm\,0.39$ $1.56\pm\,0.72$ $1.78\pm\,0.03$
[…]
$\Delta_{E}(1),\ {N_{f}=7}$ $1.64\pm\,0.47$ $1.48\pm\,0.72$ $1.78\pm\,0.03$
Table 3: Evaluations of PB-resummed $\Delta_{E}(1)$ with $N_{f}=2\sim 7$.
$N_{f}$ 22 33 44 55 66 77
$\Delta_{E}$ $1.94\pm 0.56$ $1.76\pm 0.67$ $1.64\pm 0.71$ $1.56\pm 0.72$ $1.51\pm 0.72$ $1.48\pm 0.72$
@@
Table 4: Estimates of 3D $\Delta_{E}$ obtained from the PB resummation of the four-loop truncated $(2+\epsilon)$D perturbative expansion, with $N_{f}=2\sim 7$.
S4.2 Extrapolation of $\Delta_{\phi}$
Replacing $\epsilon^{n}$ with $\delta^{n}/n!$, we obtain the Borel transform of the truncated $\Delta_{\phi}(\epsilon)$ series (S4.5):
$\mathcal{B}[\Delta_{\phi}](\delta)\ =\ \left(1-\frac{1}{N_{f}^{2}}\right)\left[\delta-\frac{\delta^{2}}{4}+\frac{\delta^{3}}{12}\right]+\mathcal{O}\left(\delta^{4}\right)\,.$ (S4.9)
If $\Delta_{\phi}(\epsilon)$ is indeed alternating, we would expect $\mathcal{B}[\Delta_{\phi}](\delta)$ to have a leading singularity on the negative real axis, and the Borel integral might be free of imaginary ambiguities. The near-diagonal Padé approximants up to order $\delta^{3}$ are
$\mathcal{B}[\Delta_{\phi}](\delta)\big|_{[1/1]}$ $=\left(1-\frac{1}{N_{f}^{2}}\right)\frac{\delta}{1+\frac{1}{4}\delta}\,$, (S4.10a)
$\mathcal{B}[\Delta_{\phi}](\delta)\big|_{[2/1]}$ $=\left(1-\frac{1}{N_{f}^{2}}\right)\frac{\delta+\frac{1}{12}\delta^{2}}{1+\frac{1}{3}\delta}\,$, (S4.10b)
$\mathcal{B}[\Delta_{\phi}](\delta)\big|_{[1/2]}$ $=\left(1-\frac{1}{N_{f}^{2}}\right)\frac{\delta}{1+\frac{1}{4}\delta-\frac{1}{48}\delta^{2}}\,.$ (S4.10c)
The $[1/1]$ and $[2/1]$ Padé approximants have no singularities on the positive real axis, while the $[1/2]$ Padé approximant has a pole at $\delta\approx 15.2\gg 0$. We then obtain the $[m/n]$ PB-resummed function $\Delta_{\phi}(\epsilon)$ by performing the Borel integral,
$\Delta_{\phi}(\epsilon)\big|_{[m/n]}=\frac{1}{\epsilon}\int_{0}^{\infty}d\delta\,e^{-\delta/\epsilon}\,\mathcal{B}[\Delta_{\phi}](\delta)\big|_{[m/n]}\,.$ (S4.11)
We show the functions $\Delta_{\phi}(\epsilon)|_{[m/n]}\big/\bigl(1-\frac{1}{N_{f}^{2}}\bigr)$ in Fig. 3. The solid lines show the real parts while the shadow bands indicate the imaginary ambiguities, which only appear for $[1/2]$ and are essentially invisible. (The imaginary ambiguities only become appreciable for $\epsilon\gtrsim 2$.) We summarize the results at $\epsilon=1$ in Table 5. We take the union of the envelopes of $\Delta_{\phi}(1)|_{[1/1]}$, $\Delta_{\phi}(1)|_{[2/1]}$, and $\Delta_{\phi}(1)|_{[2/1]}$ as our final estimates of 3D $\Delta_{\phi}$, as we summarize in Table 6. We would need more terms beyond the three-loop order in the $(2+\epsilon)$D expansion to get more accurate results.
Figure 3: Evaluations of PB-resummed $\Delta_{\phi}(\epsilon)\big/\bigl(1-\frac{1}{N_{f}^{2}}\bigr)$. The solid lines show $\mathrm{Re}\,\Delta_{\phi}(\epsilon)\big/\bigl(1-\frac{1}{N_{f}^{2}}\bigr)$ and the shadow bands indicate $\pm\mathrm{Im}\,\Delta_{\phi}(\epsilon)\big/\bigl(1-\frac{1}{N_{f}^{2}}\bigr)$. The imaginary ambiguities only appear for $[1/2]$ and are too small to be visible.
$[m/n]$ $[1/1]$ $[2/1]$ $[1/2]$
$\Delta_{\phi}(1)\big/\bigl(1-\frac{1}{N_{f}^{2}}\bigr)$ $0.70$ $0.73$ $0.74\pm i\,0.00$
Table 5: Evaluations of PB-resummed $\Delta_{\phi}(1)\big/\bigl(1-\frac{1}{N_{f}^{2}}\bigr)$.
$N_{f}$ 22 33 44 55 66 77
$\Delta_{\phi}$ $0.54\pm 0.02$ $0.64\pm 0.02$ $0.67\pm 0.02$ $0.69\pm 0.02$ $0.70\pm 0.02$ $0.71\pm 0.02$
Table 6: Estimates of 3D $\Delta_{\phi}$ obtained from the PB resummation of the three-loop truncated $(2+\epsilon)$D perturbative expansion, with $N_{f}=2\sim 7$.